# Charting the main sequence of star-forming galaxies out to redshifts $z\sim 7$

### v1

@@ 3.2 Far-IR fluxes
$\sigma^{2}(S_{\nu})=\frac{1}{\sum_{i}1/\sigma_{i}^{2}},$ (4)
@@
with the example 500 ${\rm\mu m}$ stamps for each redshift and stellar mass bin shown in Figure 9.
@@
The stacked average far-IR flux densities at each bin were then extracted from the non-linear least squares best fits to the light profiles derived from corresponding stamps. The reason for this is that, at the Herschel 250, 350 and 500 ${\rm\mu m}$ SPIRE bands, due to the relatively large beam sizes, the clustering of objects associated with the stacked sources is seen to significantly contribute to the mean flux. Figure 2 shows the light profile and the best fits for the sample of Herschel SPIRE 500 ${\rm\mu m}$ sources at $1.751.25<z<$ 1.75 with stellar masses between $10.510.25<{\rm log}(M_{\ast}/{\rm M_{\odot}})<$ 10.5. As can be seen, the stacked flux (red line) consists of the average flux from the stacked sample (blue dashed line) and the background flux from the associated sources (green dotted line), both of which can be modeled with a Gaussian function. In cases when a double Gaussian (${\rm beam+background}$) gives better fits than a single Gaussian (beam), the background is subtracted and the resulting average flux density of the stacked sample is extracted from a beam component of the best fit. The flux density errors are then derived from the corresponding covariance matrices.
@@
Figure 2: The $1.751.25<z<$ 1.75, $10.510.25<{\rm log}(M_{\ast}/{\rm M_{\odot}})<$ 10.5 light profile derived at 500 ${\rm\mu m}$. The red points represent mean values of the flux densities derived from stacked images as a function of distance from the image center. The blue dashed, green dotted and red solid lines represent the corresponding beam, the contribution from background sources and the sum of the two, respectively. For Herschel SPIRE bands, the contribution from background sources is significant and must, therefore, be subtracted from the overall light profile, when deriving stacked flux densities (see Section 3.2).
4 The SFR-stellar mass relations
4.1 SFR measurements
@@
A total star formation rate consists of the UV and IR components, where the rest-frame UV light is sensitive to the young massive stars and is therefore a good tracer of the recent, unobscured star formation, while the IR emission accounts for the stellar emission that has been absorbed by the interstellar dust, where:
${\rm SFR_{tot}}=\mathcal{K}_{\rm FUV}L_{\rm FUV}+\mathcal{K}_{\rm IR}L_{\rm IR}.$ (5)
@@
The far-UV (FUV) part of the SFR is derived from the FUV luminosity measured at 150 nm, where $L_{\rm FUV}=\nu L_{\nu}$ and $\mathcal{K}_{\rm FUV}=2.5\times 10^{-10}\,{\rm M_{\odot}\,yr^{-1}\,L_{\odot}^{-1}}$ (Kennicutt 1998), with an additional multiplicative factor of 0.63 to convert from a Salpeter to a Chabrier IMF. The FUV SFRs were calculated for each galaxy and then the mean value was evaluated at each redshift/stellar mass bin. In order to find the IR portion of the SFR we follow Madau & Dickinson (2014), where the IR luminosity is converted to the SFR using $\mathcal{K}_{\rm IR}=1.73\times 10^{-10}\,{\rm M_{\odot}\,yr^{-1}\,L_{\odot}^{-1}}$, again with an additional multiplicative factor of 0.63 to convert from a Salpeter to a Chabrier IMF.
@@
Figure 3: SFR-$M_{\ast}$ relations derived in Section 4 for the whole (star-forming and quiescent – left panel) and the star-forming (main sequence – right panel) galaxy samples. The best-fit lines, color-coded with redshift, are derived according to Equation 7.
@@
In order to find the IR luminosity, $L_{\rm IR}$, the dust emission curve of Casey (2012) was fit to the available far-IR photometry using the non-linear least squares analysis. The adopted dust emission curve, described in Table 1 of Casey (2012), can be approximated as a classic single-temperature greybody function at far-IR wavelengths, with the addition of the power-law fit at mid-IR wavelengths. The data was initially fit with both the single greybody and the power-law greybody curves and it was found that the latter gives better fits to the data at mid-IR bands. The power-law greybody curve has 4 free parameters – mid-IR power-law slope, $\alpha$, the emissivity index, $\beta$, the effective dust temperature, $T^{\rm eff}_{\rm d}$, and the overall normalization, $N$. Since, at $z\gtrsim 2$, the available photometry redwards of the dust emission peak consists only of the 850 and 500 ${\rm\mu m}$, with the 500 ${\rm\mu m}$ being affected by blending (see Figure 2), we decided to fix $\beta$ at 1.96. Similarly, due to the lack of photometry at rest-frame wavelength $\lambda\lesssim 80\,{\rm\mu m}$, we also fix $\alpha$ at 2.3. Values of $\alpha$ and $\beta$ were fixed according to the recommendations of Drew & Casey (2022). The curve is then integrated between 8 and 1000 ${\rm\mu m}$ and the errors are estimated from the corresponding covariance matrix. The stacked photometry and the best fits for the whole (star-forming and quiescent) and star-forming subsamples are shown in Figures 10 and 11, respectively. The resulting best-fit values of the total SFR at each redshift/stellar mass bin for the whole and star-forming subsamples are listed in the top and bottom panels of Table 2, as well as depicted as color points with error bars in the left and right panels of Figure 3, respectively.
4.2 Fitting the SFR-stellar mass function
@@
Since the SFR-$M_{\ast}$ relation was shown to exhibit a turnover at high masses (e.g. Lee et al. 2015; Popesso et al. 2023), we use the functional form of the SFR-$M_{\ast}$ relation from Lee et al. (2015):
@@
where the SFR transitions from a power-law with slope $\gamma$ at masses below $M_{0}$, to an asymptotic value, ${\rm SFR_{max}}$, at $M_{\ast}\gg M_{0}$. The slope, $\gamma$, has been shown to vary with redshift, with values between $0.8\lesssim\gamma\lesssim 1.3$ (e.g. Lee et al. 2015; Tomczak et al. 2016). Because of the low dynamic range in $M_{\ast}$, particularly at high redshift, we decided to fix $\gamma$. With the best-fit value of $\gamma=1.0$, adopted from Popesso et al. (2023), Equation 6 becomes:
@@
We find that the most accurate fits to the SFR-$M_{\ast}$ relation are produced when the logarithm of the free parameters of Equation 7 are assumed to follow an exponential evolution with redshift, where:
@@
Table 1: Best-fit values to parameters from Equation 8 for the star-forming and all galaxies (star-forming and quiescent).
@@
a1 $\phantom{-1}3.69\pm 0.32$ $\phantom{-1}3.47\pm 0.22$
a2 13.81±0.27plus-or-minus3.810.27-3.81$\pm$ 0.27- 3.81 ± 0.27 13.13±0.16plus-or-minus3.130.16-3.13$\pm$ 0.16- 3.13 ± 0.16
a3 $\phantom{-1}0.47\pm 0.07$ $\phantom{-1}0.56\pm 0.08$
b1 $\phantom{-}11.91\pm 0.45$ $\phantom{-}11.69\pm 0.31$
b2 12.48±0.38plus-or-minus2.480.38-2.48$\pm$ 0.38- 2.48 ± 0.38 11.66±0.24plus-or-minus1.660.24-1.66$\pm$ 0.24- 1.66 ± 0.24
b3 $\phantom{-1}0.44\pm 0.15$ $\phantom{-1}0.53\pm 0.21$
The functional form of Equation 7 is fitted to the stacked data (Table 2) using the non-linear least square fitting, where, initially, the fits are performed for the whole sample, including both star-forming and quiescent galaxies. The best-fit curves can be seen in the left panel of Figure 3 and the best-fit parameters from Equation 8 for the whole sample are listed in Table 1.
As can be seen in Figure 3, the SFR-$M_{\ast}$ relation shows a clear turnover at high stellar masses, represented in Equation 7 by the parameter $M_{0}$. However, since the whole sample consists of both the star-forming and quiescent galaxies, the apparent turnover may simply be a consequence of the increasing population of quiescent galaxies at low redshifts. We, therefore, repeat the procedure for the star-forming galaxies only. The corresponding best SFR-$M_{\ast}$ fits are shown in the right panel of Figure 3, with the best-fit parameters listed in Table 1. As can be seen, the turnover is still visible at low redshifts, albeit slightly less pronounced.
@@ 5 Discussion
Figure 4: Left panel The peak dust temperatures, $T_{\rm d}$, as a function of the IR luminosity, $L_{\rm IR}$, color-coded with redshift. It can be seen that while $T_{\rm d}$ increases with $L_{\rm IR}$ between $z\sim 0.5-6$, it does not seem to significantly vary at any given redshift. Right panel Since $T_{\rm d}$ variations with $L_{\rm IR}$ at any given redshift remain small, we plot the average peak dust temperature as a function of redshift as black points, with the best-fit black solid line. For comparison, we show the relations of Schreiber et al. (2018), Viero et al. (2022) and Ismail et al. (2023) as red dotted, green dash-dotted and blue dashed lines, respectively. For discussion see Section 5.2.
@@ 5.1 Evolution of the star-forming main sequence
As discussed in previous studies (e.g. Santini et al. 2014; Wang et al. 2022), the shape of the star-forming main sequence is controlled by the relationship between the molecular gas mass and star formation rate surface densities – the Schmidt-Kennicutt relation. After stacking a mass-complete sample of $0.4\lesssim z\lesssim 3.6$, $M_{\ast}>10^{10}\,{\rm M_{\odot}}$ main-sequence galaxies, Wang et al. (2022) found that all the MS galaxies lie on the stellar mass and redshift independent MS-only SK relation, with the slope of $\sim 1.13$. For comparison, at high $\Sigma M_{\rm gas}$, starburst galaxies exhibit two to three times higher star-forming efficiencies, producing the SK relation with slightly higher normalization (e.g. Daddi et al. 2010; Kennicutt & De Los Reyes 2021). In addition, Wang et al. (2022) showed that the FIR sizes of all the MS galaxies do not vary with stellar mass or redshift, with the mean half-light radius of $\sim 2.2$ kpc, implying the so-called integrated Schmidt-Kennicutt law, where ${\rm SFR}\propto M_{\rm gas}^{1.13}$. The star-forming main-sequence stellar mass and redshift evolution is, therefore, driven by the shape of the integrated SK relation and galaxy molecular gas content ($f_{\rm gas}$). The molecular gas fraction, $f_{\rm gas}\equiv M_{\rm gas}/M_{\ast}$, on the other hand, is driven by many processes, including the star formation efficiency (${\rm SFE\equiv SFR}/M_{\rm gas}$), gas cooling, feedback, merger rate, etc. As shown by Santini et al. (2014), the redshift and stellar mass evolution of the molecular gas content in star-forming galaxies, together with the redshift-independent SK relation, give rise to the fundamental $f_{\rm gas}$-$M_{\ast}$-SFR three-dimensional surface. Santini et al. (2014) stacked GOODS-S, GOODS-N and the COSMOS star-forming galaxies out to $z=2.5$ in Herschel bands to show that the molecular gas fraction at given stellar mass and SFR does not depend on redshift. The projection of the $f_{\rm gas}$-$M_{\ast}$-SFR surface onto the $M_{\ast}$-SFR plane, given $f_{\rm gas}$ variations with stellar mass and redshift, produces the evolution of the star-forming main sequence.
@@
From the best-fit function describing the MS evolution below the bending mass $M_{0}$ (Equations 6 and 8) at any given redshift, the SFR dependence on the stellar mass is close to linear ($\gamma\sim 1$), consistent with the constancy of the low-mass power-law slope of the stellar mass function of $\alpha\sim 1.5$ (e.g. Tomczak et al. 2014, although see Section 5.4). Since, from the MS-only integrated SK relation, ${\rm SFR}\propto M_{\rm gas}^{1.13}$, linear dependence of SFR on stellar mass implies that at given redshift and below $M_{0}$, the gas fraction should decrease with stellar mass, as was indeed found in previous works (e.g. Schreiber et al. 2016; Wang et al. 2022). Similarly, since ${\rm SFR}\propto M_{\ast}$ below $M_{0}$ and the gas fraction decreases with stellar mass, the star-forming efficiency ($\equiv{\rm SFR}/M_{\rm gas}$) will increase with $M_{\ast}$. On the other hand, for a given stellar mass, increasing the SFR with redshift (evolution of the main-sequence normalization) should also increase the molecular gas mass (MS-only SK relation). For a given stellar mass, the gas fraction must, therefore, increase with redshift, consistently with previous findings (e.g. Geach et al. 2011; Wang et al. 2022).
@@
At masses $>M_{0}$ the main sequence flattens. As suggested (e.g. Daddi et al. 2022; Popesso et al. 2023), the SFR-$M_{\ast}$ turnover is most likely caused by the phasing out of cold accretion. Based on theoretical scenarios, the cold gas can accrete freely for dark-matter (DM) halos with $M_{\rm DM}<M_{\rm shock}$, where $M_{\rm shock}$ ($\simeq 10^{11.8}\,{\rm M_{\odot}}$ independent of redshift) marks the mass above which shocks can efficiently heat incoming baryons. Above $M_{\rm shock}$ the cold accretion can still occur at higher redshifts through cold, collimated streams for halos with $M_{\rm DM}<M_{\rm stream}$, with $M_{\rm stream}$ boundary evolving with redshift from $\sim 10^{12.5}\,{\rm M_{\odot}}$ at $z=2$ to $\sim 10^{13.5}\,{\rm M_{\odot}}$ at $z=3$. Converting from halo to stellar masses using the prescription of Behroozi et al. (2013), Daddi et al. (2022) shows in their Figure 2 that the $M_{\rm shock}$-$M_{\rm stream}$ boundary is in remarkable agreement with the redshift evolution of the SFR-$M_{\ast}$ turnover mass $M_{0}$ (increasing exponentially towards higher redshifts; Equation 8). This suggests that the flattening of the SFR-$M_{\ast}$ relation, at a given redshift, is most likely caused by the suppression of the cold-gas feeding of the galaxy, which effectively decreases its molecular gas fraction. At the same time, Daddi et al. (2022) points out that this process is not equivalent to quenching, describing galaxies with very low amounts of SFR, for which additional processes are required, most likely candidates of which include mergers and AGNs. Flattening of the main sequence at $M_{\ast}>M_{0}$ gives a relatively small increase in SFR for large increase in stellar mass. Provided that main-sequence galaxies lie on the integrated SK relation, small increase in SFR gives small increase in molecular gas mass, causing the molecular gas fractions above $M_{0}$ to decrease significantly, consistent with the above scenario. However, Schreiber et al. (2016) found that flattening of the star-forming main sequence at $z\sim 1$ is not caused by the decrease of the molecular gas fraction but rather the significant drop in the star-forming efficiency, with only a small change in the $f_{\rm gas}$. For galaxy stellar masses above the bending mass, the large increase in $M_{\ast}$ gives only a small increase in SFR. A small change in $f_{\rm gas}$ would imply that the corresponding molecular gas mass should also significantly increase. This is inconsistent with the finding of Wang et al. (2022), that all the star-forming galaxies lie on the SK relation, where ${\rm SFR}\propto M_{\rm gas}^{1.13}$, since small increase in SFR above $M_{0}$ for main-sequence galaxies should give similarly small rise in the molecular gas mass (and hence large decrease in $f_{\rm gas}$). A near-constant gas fraction above $M_{0}$ would, therefore, suggest that the main-sequence galaxies at these large masses should lie below the SK relation. More work, with larger samples, is clearly required to resolve this inconsistency.

### v2

@@ 3.2 Far-infrared fluxes
$\sigma_{\rm stack}^{2}(S_{\nu})=\frac{1}{\sum_{i}1/\sigma_{i}^{2}},$ (4)
@@
with the example 500 ${\rm\mu m}$ stamps for each redshift and stellar mass bin given in Sect. 7.
@@
The stacked average FIR flux densities at each bin were then extracted from the nonlinear least squares best fits to the light profiles derived from corresponding stamps. The reason for this is that, at the Herschel 250, 350, and 500 ${\rm\mu m}$ SPIRE bands, due to the relatively large beam sizes, the clustering of objects associated with the stacked sources is seen to significantly contribute to the mean flux. Figure 2 shows the light profile and the best fits for the sample of Herschel SPIRE 500 ${\rm\mu m}$ sources at $1.751.25<z<$ 1.75 with stellar masses between $10.510.25<{\rm log}(M_{\ast}/{\rm M_{\odot}})<$ 10.5. As can be seen, the stacked flux (red line) consists of the average flux from the stacked sample (dashed blue line) and the background flux from the associated sources (dotted green line), both of which can be modeled with a Gaussian function. In cases when a double Gaussian (${\rm beam+background}$) gives better fits than a single Gaussian (beam), the background is subtracted and the resulting average flux density of the stacked sample is extracted from a beam component of the best fit.
The errors on the stacked FIR fluxes, which are related to the redshifts, stellar masses, and the $NUVrJ$ quiescent galaxy selection procedure uncertainties, were estimated using the bootstrapping method. In each step, a mock catalog was constructed by drawing, at random and with replacement, a sample of sources from the original mass-complete dataset. The stacking procedure, described above, was then applied and the average fluxes were extracted. The process was repeated 1000 times and the uncertainties on the FIR fluxes, $\sigma_{\rm bootstrap}$, taken to be the standard deviation of the resulting simulated values. In addition, the contribution to the final stacked FIR fluxes errors from the uncertainties related to the functional fits to the light profiles, $\sigma_{\rm fit}$, described above, as well as the errors related to the individual observed FIR fluxes from Eq. 4 were included, with the final errors on the stacked values equal to
$\sigma(S_{\nu})=\sqrt{\sigma_{\rm bootstrap}^{2}+\sigma_{\rm fit}^{2}+\sigma_{\rm stack}^{2}}.$ (5)
@@
Figure 2: $1.751.25<z<$ 1.75, $10.510.25<{\rm log}(M_{\ast}/{\rm M_{\odot}})<$ 10.5 light profile derived at 500 ${\rm\mu m}$. The red points represent mean values of the flux densities derived from stacked images as a function of distance from the image center. The dashed blue, dotted green, and solid red lines represent the corresponding beam, the contribution from background sources, and the sum of the two, respectively. For Herschel SPIRE bands, the contribution from background sources is significant and must therefore be subtracted from the overall light profile, when deriving stacked flux densities (see Sect. 3.2).
4 The star formation rate–stellar mass relations
4.1 Star formation rate measurements
@@
A total SFR consists of the UV and IR components, where the rest-frame UV light is sensitive to the young massive stars and is therefore a good tracer of recent, unobscured star formation, while the IR emission accounts for the stellar emission that has been absorbed by the interstellar dust. Here, we followed the recent calibration of Kennicutt & Evans (2012), where
${\rm SFR_{tot}}=1.71\times 10^{-10}(L_{\rm FUV}+0.46\times L_{\rm IR}).$ (6)
@@
The far-UV (FUV) part of the SFR was derived from the FUV luminosity measured at 150 nm, where $L_{\rm FUV}=\nu L_{\nu}$, with FUV SFRs calculated for each galaxy and then the mean value evaluated at each redshift or stellar mass bin.
@@
Figure 3: SFR-$M_{\ast}$ relations derived in Sect. 4 for the whole (star-forming and quiescent – left panel) and the star-forming (MS – right panel) galaxy samples. The best-fit lines, color-coded with redshift, were derived according to Eq. 8.
@@
In order to find the IR luminosity, $L_{\rm IR}$, the dust emission curve of Casey (2012) was fit to the available FIR photometry using the nonlinear least squares analysis. The adopted dust emission curve, described in Table 1 of Casey (2012), can be approximated as a classic single-temperature gray-body function at FIR wavelengths, with the addition of the power law fit at mid-IR wavelengths. The data was initially fit with both the single gray-body and the power law gray-body curves and it was found that the latter gives better fits to the data at mid-IR bands. The power law gray-body curve has four free parameters – the mid-IR power law slope, $\alpha$, the emissivity index, $\beta$, the effective dust temperature, $T^{\rm eff}_{\rm d}$, and the overall normalization, $N$. Since, at $z\gtrsim 2$, the available photometry redward of the dust emission peak consists only of the 850 and 500 ${\rm\mu m}$, with the 500 ${\rm\mu m}$ being affected by blending (see Fig. 2), we decided to fix $\beta$ at 1.96. Similarly, due to the lack of photometry at a rest-frame wavelength of $\lambda\lesssim 80\,{\rm\mu m}$, we also fixed $\alpha$ at 2.3. Values of $\alpha$ and $\beta$ were fixed according to the recommendations of Drew & Casey (2022). The curve was then integrated between 8 and 1000 ${\rm\mu m}$ and the errors were estimated from the corresponding covariance matrix. The stacked photometry and the best fits for the whole (star-forming and quiescent) and star-forming subsamples are given in Sect. 7. The resulting best-fit values of the total SFR at each redshift or stellar mass bin for the whole and star-forming subsamples are listed in the top and bottom panels of Table 2, as well as being depicted as color points with error bars in the left and right panels of Fig. 3, respectively.
4.2 Fitting the star formation rate–stellar mass function
@@
Since the SFR-$M_{\ast}$ relation was shown to exhibit a turnover at high masses (e.g., Lee et al. 2015; Popesso et al. 2023), we used the functional form of the SFR-$M_{\ast}$ relation from Lee et al. (2015):
@@
where the SFR transitions from a power law with slope $\gamma$ at masses below $M_{0}$ to an asymptotic value, ${\rm SFR_{max}}$, at $M_{\ast}\gg M_{0}$. The slope, $\gamma$, has been shown to vary with redshift, with values between $0.8\lesssim\gamma\lesssim 1.3$ (e.g., Lee et al. 2015; Tomczak et al. 2016). Because of the low dynamic range in $M_{\ast}$, particularly at high redshift, we decided to fix $\gamma$. With the best-fit value of $\gamma=1.0$, adopted from Popesso et al. (2023), Eq. 7 becomes
@@
We find that the most accurate fits to the SFR-$M_{\ast}$ relation are produced when the logarithm of the free parameters of Eq. 8 are assumed to follow an exponential evolution with redshift, where
@@
Table 1: Best-fit values to parameters from Eq. 9 for the star-forming galaxies and for all galaxies (both star-forming and quiescent).
@@
a1 $\phantom{-1}3.28\pm 0.32$ $\phantom{-1}2.97\pm 0.20$
a2 13.36±0.26plus-or-minus3.360.26-3.36$\pm$ 0.26- 3.36 ± 0.26 12.62±0.14plus-or-minus2.620.14-2.62$\pm$ 0.14- 2.62 ± 0.14
a3 $\phantom{-1}0.46\pm 0.08$ $\phantom{-1}0.59\pm 0.10$
b1 $\phantom{-}11.73\pm 0.73$ $\phantom{-}11.38\pm 0.39$
b2 12.06±0.62plus-or-minus2.060.62-2.06$\pm$ 0.62- 2.06 ± 0.62 11.14±0.30plus-or-minus1.140.30-1.14$\pm$ 0.30- 1.14 ± 0.30
b3 $\phantom{-1}0.37\pm 0.23$ $\phantom{-1}0.50\pm 0.36$
The functional form of Eq. 8 was fit to the stacked data (Table 2) using the nonlinear least square fitting, where, initially, the fits were performed for the whole sample, including both star-forming and quiescent galaxies. The best-fit curves can be seen in the left panel of Fig. 3 and the best-fit parameters from Eq. 9 for the whole sample are listed in Table 1.
As can be seen in Fig. 3, the SFR-$M_{\ast}$ relation shows a clear turnover at high stellar masses, represented in Eq. 8 by the parameter $M_{0}$. However, since the whole sample consists of both the star-forming and quiescent galaxies, the apparent turnover may simply be a consequence of the increasing population of quiescent galaxies at low redshifts. We therefore repeated the procedure for the SFGs only. The corresponding best SFR-$M_{\ast}$ fits are shown in the right panel of Fig. 3, with the best-fit parameters listed in Table 1. As can be seen, the turnover is still visible at low redshifts, albeit slightly less pronounced.
@@ 5 Discussion
Figure 4: $M_{\rm shock}$ and $M_{\rm stream}$ boundaries, above which the star formation is theoretically expected to be quenched, as is described in Sect. 5.1, shown in blue, where halo masses were converted to stellar masses using the prescription of Behroozi et al. (2013). The empirical curve of Behroozi et al. (2013) is also shown with a solid green line. For comparison, the results of this work are plotted in black. Red data represents the $M_{0}$ redshift evolution found in Popesso et al. (2023).
@@ 5.1 Evolution of the star-forming main sequence
As has been discussed in previous studies (e.g., Santini et al. 2014; Wang et al. 2022), the shape of the star-forming MS is controlled by the relationship between the molecular gas mass and SFR surface densities – the SK relation. After stacking a mass-complete sample of $0.4\lesssim z\lesssim 3.6$, $M_{\ast}>10^{10}\,{\rm M_{\odot}}$ MS galaxies, Wang et al. (2022) found that all the MS galaxies lie on the stellar mass and redshift independent MS-only SK relation, with the slope of $\sim 1.13$. For comparison, at high $\Sigma M_{\rm gas}$, starburst galaxies exhibit two to three times higher star-forming efficiencies, producing the SK relation with slightly higher normalization (e.g., Daddi et al. 2010; Kennicutt & De Los Reyes 2021). In addition, Wang et al. (2022) showed that the FIR sizes of all the MS galaxies do not vary with stellar mass or redshift, with the mean half-light radius of $\sim 2.2$ kpc, implying the so-called integrated SK law, where ${\rm SFR}\propto M_{\rm gas}^{1.13}$. The star-forming MS stellar mass and redshift evolution is therefore driven by the shape of the integrated SK relation and galaxy molecular gas content ($f_{\rm gas}$). The molecular gas fraction, $f_{\rm gas}\equiv M_{\rm gas}/M_{\ast}$, on the other hand, is driven by many processes, including the star formation efficiency (${\rm SFE\equiv SFR}/M_{\rm gas}$), gas cooling, feedback, and the merger rate. As has been shown by Santini et al. (2014), the redshift and stellar mass evolution of the molecular gas content in SFGs, together with the redshift-independent SK relation, give rise to the fundamental $f_{\rm gas}$-$M_{\ast}$-SFR three-dimensional surface. Santini et al. (2014) stacked GOODS-S, GOODS-N and the COSMOS SFGs out to $z=2.5$ in Herschel bands to show that the molecular gas fraction at given stellar mass and SFR does not depend on redshift. The projection of the $f_{\rm gas}$-$M_{\ast}$-SFR surface onto the $M_{\ast}$-SFR plane, given $f_{\rm gas}$ variations with stellar mass and redshift, produces the evolution of the star-forming MS.
@@
From the best-fit function describing the MS evolution below the bending mass $M_{0}$ (Eqs. 7 and 9) at any given redshift, the SFR dependence on the stellar mass is close to linear ($\gamma\sim 1$), consistent with the constancy of the low-mass power law slope of the stellar mass function of $\alpha\sim 1.5$ (e.g., Tomczak et al. 2014, although see Sect. 5.4). Since, from the MS-only integrated SK relation, ${\rm SFR}\propto M_{\rm gas}^{1.13}$, a linear dependence of the SFR on stellar mass implies that at a given redshift and below $M_{0}$, the gas fraction should decrease with stellar mass, as was indeed found in previous works (e.g., Schreiber et al. 2016; Wang et al. 2022). Similarly, since ${\rm SFR}\propto M_{\ast}$ below $M_{0}$ and the gas fraction decreases with stellar mass, the star-forming efficiency ($\equiv{\rm SFR}/M_{\rm gas}$) will increase with $M_{\ast}$. On the other hand, for a given stellar mass, increasing the SFR with redshift (evolution of the MS normalization) should also increase the molecular gas mass (MS-only SK relation). For a given stellar mass, the gas fraction must therefore increase with redshift, which is consistent with previous findings (e.g., Geach et al. 2011; Wang et al. 2022).
@@
At masses $>M_{0}$, the MS flattens. As has been suggested (e.g., Daddi et al. 2022; Popesso et al. 2023), the SFR-$M_{\ast}$ turnover is most likely caused by the phasing out of cold accretion. Based on theoretical scenarios, the cold gas can accrete freely for dark matter halos with $M_{\rm DM}<M_{\rm shock}$, where $M_{\rm shock}$ ($\simeq 10^{11.8}\,{\rm M_{\odot}}$ independent of redshift) marks the mass above which shocks can efficiently heat incoming baryons. Above $M_{\rm shock}$, the cold accretion can still occur at higher redshifts through cold, collimated streams for halos with $M_{\rm DM}<M_{\rm stream}$, with $M_{\rm stream}$ boundary evolving with redshift from $\sim 10^{12.5}\,{\rm M_{\odot}}$ at $z=2$ to $\sim 10^{13.5}\,{\rm M_{\odot}}$ at $z=3$. Converting from halo to stellar masses using the prescription of Behroozi et al. (2013), we plot the redshift evolution of the $M_{\rm shock}$-$M_{\rm stream}$ boundary in Fig. 4, with the $M_{\rm shock}$, $M_{\rm stream}$ and $M_{\rm shock}+M_{\rm stream}$ boundaries shown in dotted, dash-dotted, and solid blue lines, respectively. We also plot the results of this work in black, where it can be seen that the boundary is in good agreement with the redshift evolution of the SFR-$M_{\ast}$ turnover mass $M_{0}$ (increasing exponentially toward higher redshifts; Eq. 9). For comparison, the results of Behroozi et al. (2013) and Popesso et al. (2023) are also shown. This suggests that the flattening of the SFR-$M_{\ast}$ relation, at a given redshift, is most likely caused by the suppression of the cold-gas feeding of the galaxy, which effectively decreases its molecular gas fraction. At the same time, Daddi et al. (2022) points out that this process is not equivalent to quenching, describing galaxies with very low SFRs, for which additional processes are required, the most likely candidates of which include mergers and AGNs.
Flattening of the MS at $M_{\ast}>M_{0}$ gives a relatively small increase in SFR for a large increase in stellar mass. Provided that MS galaxies lie on the integrated SK relation, a small increase in SFR gives a small increase in molecular gas mass, causing the molecular gas fractions above $M_{0}$ to decrease significantly, consistent with the above scenario. However, Schreiber et al. (2016) found that flattening of the star-forming MS at $z\sim 1$ is caused not by the decrease in the molecular gas fraction but rather by the significant drop in the star-forming efficiency, with only a small change in $f_{\rm gas}$. For galaxy stellar masses above the bending mass, the large increase in $M_{\ast}$ gives only a small increase in SFR. A small change in $f_{\rm gas}$ would imply that the corresponding molecular gas mass should also significantly increase. This is inconsistent with the finding of Wang et al. (2022) that all the SFGs lie on the SK relation, where ${\rm SFR}\propto M_{\rm gas}^{1.13}$, since a small increase in the SFR above $M_{0}$ for MS galaxies should give similarly small rise in the molecular gas mass (and hence large decrease in $f_{\rm gas}$). A near-constant gas fraction above $M_{0}$ would therefore suggest that the MS galaxies at these large masses should lie below the SK relation. More work, with larger samples, is clearly required to resolve this inconsistency.
Figure 5: Evolution of peak dust temperature found in this work. Left panel: The peak dust temperatures, $T_{\rm d}$, as a function of the IR luminosity, $L_{\rm IR}$, color-coded with redshift. It can be seen that while $T_{\rm d}$ increases with $L_{\rm IR}$ between $z=0.45-5.7$, it does not seem to significantly vary at any given redshift. Right panel: Since $T_{\rm d}$ variations with $L_{\rm IR}$ at any given redshift remain small, we plot the average peak dust temperature as a function of redshift as black points, with the best-fit solid black line. For comparison, we show the relations of Schreiber et al. (2018), Viero et al. (2022) and Ismail et al. (2023) as dotted red, dash-dotted green, and dashed blue lines, respectively. For discussion see Sect. 5.2.